# The Anderson transition – a view from Krylov space

### v3

@@ III.2 Critical Disorder
The Krylov space properties of the critical point can be constrained by the same analysis we’ve conducted on the delocalized side. The Anderson critical point is known to exhibit multi-fractality [Janssen1994_Multifractal_localization, Evers_2008_AndersonTransitions] and was recently studied in detail in Ref. [Hopjan2023_Scale_InvariantSurvivalProb]. This multi-fractality manifests itself as powerlaw scaling of the survival probability $C(t)\sim t^{-\beta}$, with $\beta\approx 0.42$. At the critical point, the Thouless time approaches the Heisenberg time (and is $O(1/V)$) implying
@@ III.3 Strong Disorder
Figure 7: The zero-mode probability density in Krylov space for Anderson models in different spatial dimensions $d$. In all cases the disorder is chosen to be much greater than the critical disorder such that all states are well localized (note the critical disorder in four dimensions is estimated to be $W_{c}\approx 33.2$ [PMarkos_1994_Metal_InsulatorTransition4D]). The results are shown as a function of the anticipated stretched Krylov coordinate $n^{1/2d}$, rescaled by the maximum value $n_{\text{max}}^{\frac{1}{2d}}$ so that all curves appear in the same scale. $d=1$ appears in the inset which has the same axes but an extended vertical range. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement. Results for spatial dimensions $d=1,2,3,4$ have disorders $W=10,20,40,80$, lattice lengths $L=12000,110,23,10$, and number of samples $N_{s}=500,1000,1000,1000$ respectively.
@@
This results in a stretched exponential zero-mode with a probability density given by $|Q_{n}|^{2}\sim\exp{(-cn^{1/2d})}$. This behavior is consistent with our numerical data for $W>16$, as shown in Fig. 6, and is additionally confirmed by our fit of the dimerization model in Fig. 3 where $\gamma=1/2d\approx 0.167$ is found consistently in the localized regime. To further substantiate our proposed scaling, we repeat the same analysis for Anderson models ranging from one to four spatial dimensions. The results shown in Fig. 7 all indicate perfect stretched exponential decay with the predicted exponent $\gamma$.
@@ IV Conclusions and Outlook
In summary, we have described the form of LIOMs in the Anderson model in Krylov space as generated by projectors on single sites. There are three types of behaviors – stretched exponential decay inside the localized phase (with geometric exponent $\gamma=1/2d$ set by real space dimension $d$), delocalization on real-space volume $V=\sqrt{D_{K}}$ states in the metallic/extended phase, and powerlaw decay at the critical point. Importantly, these results may be interpreted using effective disorder-free chains obtained from disorder averaged Krylov coefficients, which allows considerable further insight into the Anderson transition: e.g. the significance of Krylov dimerization as a mechanism of localizing operators despite having a gapless Liouvillian. The critical point occurs when the average dimerization decreases faster than a powerlaw with power 1, which is the point at which the zero-modes cease to be normalizable in Krylov space ($\gamma=0$ in expression (10)).
Our work may be extended to other aspects of the Anderson transition: for example, to perhaps enable the tracking of the mobility edge as a function of typical energy. It also sets up a framework for extending these results to interacting models where the coefficients grow with some power $b_{n}\sim n^{\delta}$ (with $\delta\leq 1$). If the dimerization decays like a powerlaw – since only the ratio of Lanczos coefficients matters for the zero-mode – the growth simply re-normalizes that power. Consequently, the faster the coefficients grow, i.e. the faster the operator growth, the slower the dimerization must decay in order for conserved quantities to exist. For spin systems debated to show MBL, such as the disordered Heisenberg model, the coefficients have been found to grow as $\sim n/\log{n}$ on average, with some dimerization [Trigueros2022_KrylovComplexityofMBL, Weisse2025_OperatorGrowthMBL]. The $\log$ correction would only be relevant at the critical power implying that the dimerization has to be constant for conserved quantities to exist. Interestingly, this still leaves the possibility for the dimerization to decay in interacting integrable models such as the Heisenberg model, where there is typically square root growth [Parker2019_universaloperatorgrowth].

### v4

@@ III.2 Critical Disorder
The Krylov space properties of the critical point can be constrained by the same analysis we’ve conducted on the delocalized side. The Anderson critical point is known to exhibit multi-fractality [36, 25] and was recently studied in detail in Ref. [31]. This multi-fractality manifests itself as powerlaw scaling of the survival probability $C(t)\sim t^{-\beta}$, with $\beta\approx 0.42$. At the critical point, the Thouless time approaches the Heisenberg time (and is $O(V)$) implying
@@ III.3 Strong Disorder
This results in a stretched exponential zero-mode with a probability density given by $|Q_{n}|^{2}\sim\exp{(-cn^{1/2d})}$. This behavior is consistent with our numerical data for $W>16$, as shown in Fig. 6, and is additionally confirmed by our fit of the dimerization model in Fig. 3 where $\gamma=1/2d\approx 0.167$ is found consistently in the localized regime. To further substantiate our proposed scaling, we repeat the same analysis for Anderson models ranging from one to four spatial dimensions. The results shown in Fig. 7 all indicate perfect stretched exponential decay with the predicted exponent $\gamma$ (Lanczos coefficients for these results are presented in Fig. 11 in Appendix E.
Figure 7: The zero-mode probability density in Krylov space for Anderson models in different spatial dimensions $d$. In all cases the disorder is chosen to be much greater than the critical disorder such that all states are well localized. The results are shown as a function of the anticipated stretched Krylov coordinate $n^{1/2d}$, rescaled by the maximum value $n_{\text{max}}^{\frac{1}{2d}}$ so that all curves appear in the same scale. $d=1$ appears in the inset which has the same axes but an extended vertical range. The dot-dashed line represents the geometric average of the zero-modes constructed in each sample whereas the colored lines are obtained from the sample-averaged $b_{n}$: the two are in complete agreement. Results for spatial dimensions $d=1,2,3,4$ have disorders $W=10,20,40,80$, lattice lengths $L=12000,110,23,10$, and number of samples $N_{s}=500,1000,1000,1000$ respectively.
@@ IV Conclusions and Outlook
In summary, we have described the form of LIOMs in the Anderson model in Krylov space as generated by projectors on single sites. There are three types of behaviors – stretched exponential decay inside the localized phase (with geometric exponent $\gamma=1/2d$ set by real space dimension $d$), delocalization on real-space volume $V=\sqrt{D_{K}}$ states in the metallic/extended phase, and powerlaw decay at the critical point. Importantly, these results may be interpreted using effective disorder-free chains obtained from disorder averaged Krylov coefficients, which allows considerable further insight into the Anderson transition: e.g. the significance of Krylov dimerization as a mechanism of localizing operators despite having a gapless Liouvillian. The critical point occurs when the average dimerization decreases faster than a powerlaw with power 1, which is the point at which the zero-modes cease to be normalizable in Krylov space ($\gamma=0$ in expression (10)). The dimerization is thus a useful probe to discriminate between ergodic and non-ergodic phases when the average operator growth is otherwise the same.
Our work may be extended to other aspects of the Anderson transition: for example, to perhaps enable the tracking of the mobility edge as a function of typical energy. It also sets up a framework for extending these results to interacting models where the coefficients grow with some power $b_{n}\sim n^{\delta}$ (with $\delta\leq 1$). If the dimerization decays like a powerlaw – since only the ratio of Lanczos coefficients matters for the zero-mode – the growth simply re-normalizes that power. Consequently, the faster the coefficients grow, i.e. the faster the operator growth, the slower the dimerization must decay in order for conserved quantities to exist. For spin systems debated to show MBL, such as the disordered Heisenberg model, the coefficients have been found to grow as $\sim n/\log{n}$ on average, with some dimerization [89, 93]. The $\log$ correction would only be relevant at the critical power implying that the dimerization has to be constant for conserved quantities to exist. Interestingly, this still leaves the possibility for the dimerization to decay in interacting integrable models such as the Heisenberg model, where there is typically square root growth [57].
@@
Appendix A Solving for the zero-modes in Krylov space
We will briefly illuminate the derivation of expression (7), which is found by solving $L_{n,m}Q_{n}=0$ for $Q_{n}$ in Krylov space:
$L_{n,m}Q_{n}=\begin{pmatrix}0&b_{2}&0&0&\cdots\\ b_{2}&0&b_{3}&0&\cdots\\ 0&b_{3}&0&b_{4}\\ 0&0&b_{4}&0&\ddots\\ \vdots&\vdots&&\ddots&\ddots\end{pmatrix}\begin{pmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\\ Q_{4}\\ \vdots\end{pmatrix}=\begin{pmatrix}0\\ 0\\ 0\\ 0\\ \vdots\end{pmatrix}$ (20)
Then, one can directly compute:
$Q_{2}b_{2}$ $=0\quad\&\quad Q_{3}=-\frac{b_{2}}{b_{3}}Q_{1}$
$Q_{4}b_{4}$ $=0\quad\&\quad Q_{5}=-\frac{b_{4}}{b_{5}}Q_{3}$
$\vdots\quad\quad\quad\quad\quad\quad\vdots$
$Q_{2n}$ $=0\quad\&\quad Q_{2n+1}=(-1)^{n}\prod_{j=1}^{n}\frac{b_{2j}}{b_{2j+1}}Q_{1},$
arriving at expression (7) for the Krylov space zero-mode.
Appendix B Variance oscillations in the extended regime variances
Figure 8: Normalized and rescaled variance of the Lanczos coefficients in the extended regime ($W=4$) for system sizes $L=14,18,22$ (top to bottom). The oscillations are due to finite size recurrences, which is shown by their collapse when rescaling the Krylov space by the linear system size $(n\rightarrow n/L)$. The vertical axis is rescaled by the volume $(\times L^{3})$ to better show the collapse.
In Fig. 1 there are oscillations in the extended regime due to finite size recurrences due to the small linear system sizes. This is clearly seen in Fig. 8 where the Krylov space has been rescaled by $n\rightarrow n/L$, showing a collapse of the oscillatory frequencies.
Appendix C A closer look at the dimerization
In Fig. 9 we show the Lanczos coefficients for a few disorders, zoomed in so the dimerized pattern is more clear. This pattern is very stable (i.e. even coefficients are larger than surrounding odd ones), and one can see by eye a reduction in the dimerization across the critical point.
Figure 9: Lanczos coefficients for $L=22$ in three dimensions and disorders (from top to bottom) $W=20,18,16,14,12$
Appendix D The dimerization fitting procedure
The parameters in Fig. 3 were found by fitting the sample-log-averaged dimerization $(\frac{b_{2n+1}}{b_{2n}}-1)$ for $n>40$ with a least squares routine following the model defined in expression (10). The dimerization was constructed in each sample first before taking the average. Within the fit, $\epsilon$ was set to be strictly non-negative, while $\gamma$ and $\alpha$ were unrestricted. As a proof of concept of this fit, we used the fitted parameters to predict the odd sample-averaged Lanczos coefficients $b_{2n+1}$ from the even ones $b_{2n}$, and plot the subsequent relative error in Fig. 10, which is shown to be both less than $10^{-2}$ across Krylov space for all disorders.
Figure 10: Relative errors for the odd sample-averaged Lanczos coefficients shown in Fig. 2 using the fitted parameters shown in Fig. 3.
Appendix E Lanczos coefficients in all spatial dimensions
Here in Fig. 11 we show the Lanczos coefficients for spatial dimensions $d=1,2,3,4$ as used to construct the zero-modes in the respective localized phases in Fig. 7.
Figure 11: Lanczos coefficients for spatial dimensions $d=1,2,3,4$ (top to bottom) with respective disorders $W=10,20,40,80$, lattice lengths $L=12000,110,23,10$, and number of samples $N_{s}=500,1000,1000,1000$ respectively.